Red light of wavelength 633 nm from a helium-neon laser passes through a slit 0.405 mm wide. The diffraction pattern is observed on a screen 3.30 m away. Define the width of a bright fringe as the distance between the minima on either side.

(a) What is the width of the central bright fringe?
mm

(b) What is the width of the first bright fringe on either side of the central one?
mm
Answer #1

Wavelength of the light, \lambda= 633nm

Width of the slits, a= 0.405mm

Distance of slits from Scree, d= 3.30m

a). The first minimum is located at

  sin \theta_1= \frac{\lambda}{a}= \frac{633 \times 10^{-9}}{0.405 \times 10^{-3}}

  \theta_1= 0.0895^\circ

Then distance of central bright fringe on one side of center is

  y_1=d tan \theta_1= 3.30tan (0.0895^\circ)= 5.15 \times 10^{-3}m= 5.15mm

Then width of central bright fringe on both sides of center will be

w_o= 2y_1=2(5.15 \times 10^{-3})m= 10.3 \times 10^{-3}m= 10.3mm(ANS)

b). Width of 1st bright fringe on either side of central bright fringe is

  w_1= y_2-y_1

where y2 is calculated using the relation,

  sin \theta_2= \frac{2 \lambda}{a}= \frac{2(633 \times 10^{-9})}{0.405 \times 10^{-3}}

\theta_2= 0.179^\circ

thus, y_2= d tan\theta_2= 3.30tan(0.179^\circ)= 0.01030m= 10.30mm

Then width of 1st bright fringe on either side of central bright fringe will be

w_1= y_2-y_1= (10.30-5.15)mm= 5.15 mm= 5.15 \times 10^{-3}m(ANS)